SpaceRef

SpaceRef


SOHO Daily Meeting Minutes for Wednesday 25 June 2003

Status Report From: Goddard Space Flight Center
Posted: Wednesday, June 25, 2003

ANNOUNCEMENTS

* We expect to lose the high gain antenna link on Wednesday or Thursday of this week for ~20 days. Starting Wednesday, we are scheduling daily passes on 34 meter stations for monitoring and commanding of the spacecraft. The 34 meter schedule for this week is included below. Since the times are very limited, we'd like to request that ALL TEAMS BE AVAILABLE (at EOF or by phone) during the passes, in case your instrument comes up in an unexpected state. When downlink only 34 meter stations are scheduled, a 26 meter station will be used for uplink (i.e. all 34 meter passes will have commanding capability).

Week 26 34-Meter Passes

                   BOT   EOT
                 --------------
   Wed     DSS-65 15:05-18:25 UT D/L only and DSS-16 26-meter overlap        
   Thurs   DSS-54 15:30-17:45 UT
   Fri     DSS-54 13:00-17:45 UT
   Sat     DSS-24 17:35-20:35 UT
   Sun     DSS-54 13:00-17:30 UT

* Current RHESSI default target is AR10386. When selecting this target to maximize coordination w/others, use campaign number 6850.

* GITT scans for the EOF and EAF are scheduled for Thursday 26 June at 9:30am.

* SPWG on Friday in the EOF at 10am.

* CNE network outage from 6-8pm on Wednesday 2 July. Internet access, e-mail services and building to building traffic will be intermittent.

FOT REPORT

Spacecraft Status: Nominal, except for the HGA pointing mechanism

HGA Position Monitoring Disabled

HGA Nominal Control Disabled

Spacecraft Anomalies:

176/0024 TT69 violated YL limits. Thermal circuit 22 duty cycle adjusted to 70% per OCD #1804.

Accomplished Activities: RSL, New tracking star, mag = 2.9, Safed CDS, UVCS, and SWAN for LR Ops, Set TC Watchdog to 72 hours per OCD #1801

Planned Activities: NONCOHO_03176

Upcoming Operations: None

Ground Anomalies:

175/1806 NRT commanding delayed 3 minutes 41 seconds due to a corrupted command at D16, cause unknown.

175/1829 SVM commanding delayed 3 minutes 13 seconds due to a corrupted command at D16, cause unknown.

175/1900 NRT commanding delayed 2 minutes 29 seconds due to a corrupted command at D16, cause unknown.

SOLAR STATUS

06/24 DOY 175 - 06/25 DOY 176

EIT daily On-Line Solar Status and observations: http://umbra.nascom.nasa.gov/eit/plan/log_form20030625.html

LASCO daily On-Line Solar Status and observations: http://lasco-www.nrl.navy.mil/observations/Todays_obs.html

Web Page for Planning: http://www.bbso.njit.edu/arm

Magnetic Maps and Heliospheric Forecasts: http://www.lmsal.com/forecast/

INSTRUMENT STATUS

CDS: Safed - doors closed.

UVCS: Safed - door closed.

LASCO: Safed - doors closed.

EIT: Bakeout.

MDI: Nominal. Will safe when we get to low rate.

TRACE: Nominal. Observing AR100386 then AR10387 & 10393. Full disk mosaics will start tomorrow.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.